# A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models

### v1

@@ Introduction
Now, we use the all-term intervention method to localize mover heads that carry term-related information. In the residual stream patching results (Figure 4(d)), we observe the highest positive score at the $[p]$ position, while negative scores are most prominent at the $[m_{1}]$ and $[m_{2}]$ positions, indicating that the information residing in these positions indeed contributes to their corresponding token prediction. This observation aligns with the importance of token embedding information at their respective positions given iterative refinement on the residual stream (Simoulin and Crabbé 2021). A closer examination of the last token position reveals that the negative effect propagates from relatively early layers (approximately layer 10 onwards), yet positive effects are from later layers (approximately layer 14 onwards), incurring a positive shift of the model’s prediction from $m$ to $p$.
@@
Given the findings that the information for prediction resides in term-specific positions, we trace which attention heads transfer the information from each term position ($[p]$, $[m_{1}]$ and $[m_{2}]$) to the last token. We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al. 2023; García-Carrasco, Maté, and Carlos Trujillo 2024). To discover the source of information flow, we investigate the heads with high attention value weight ($W_{V}$) patching score and identify 9 notable mover heads (see Appendix G for details of mover heads localization). By further investigation of attention patterns (Figure 4(f)–(h)), $\mathbf{h}_{14.14},\mathbf{h}_{15.14}$ and $\mathbf{h}_{18.12}$ are classified as positive copy heads attending highly to $[p]$, and $\mathbf{h}_{19.1}$ is a positive suppression head by attending to $[m_{2}]$ and $[s]$. $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$ and $\mathbf{h}_{23.10}$ are negative copy heads giving a strong attention to non-$[p]$.
@@
Summary
Overall, we found that the internal mechanisms to arrive at the correct conclusion involve the following phases:
  1. 1.
    Long Induction: Early layers exhibit biases towards the wrong conclusion due to long-range repetition of the first premise “All A are B”.
  2. 2.
    Duplication: Induction heads aggregate information about duplicated middle terms in the premises.
  3. 3.
    Suppression: The model aggregates and inhibits middle-term information (i.e.,“B”) suppressing the long induction mechanism.
  4. 4.
    Mover: Token-specific information is propagated to the last token position. The process culminates in the prediction shift from B’ to the correct token ‘C’.
Therefore, the discovered circuit is mainly characterized by an internal error correction mechanism. Interestingly, this mechanism is different from the way humans would reason on syllogistic arguments through the use of abstract logical primitives and inference rules.
Circuit Evaluation
@@
We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 6(a)) and belief-inconsistent (Figure 6(b)) (See Appendix H for the intervention results). Notably, the symbolic circuit proves necessary for both types of non-symbolic inputs, suggesting that the logic derived from symbolic syllogisms remains essential even when natural words are substituted. Regarding sufficiency, while belief-consistent data show significant performance recovery, we observe an inability to restore original performance on the belief-inconsistent set despite an increasing trend. These results indicate that belief biases encoded in different attention heads may play an important role.
@@
Generalization to Different Syllogistic Schemes
@@
We further aim to understand whether the symbolic circuit is specific to the AAA-1 syllogism (Barbara). To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and I for details about the syllogistic schemes and more experimental results). To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (2) Sufficiency, and (C3) Positive Logit Difference. We also measure the accuracy of the completion task for each syllogism. From Table 2, we observe that the circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%), supporting the conclusion that the circuit includes components that are crucial for the emergence of syllogistic reasoning in general. Notably, all three conditions are satisfied for the affirmative syllogisms (AII-3, IAI-3, IAI-4), for which the model achieves an accuracy above 0.6.
Generalization to Different Model Sizes
@@
Finally, we expand our analysis to different sizes of GPT-2 (small, large and XL), assuming core mechanisms would transfer across models trained on the same dataset. Here, the intervention results show similar suppression mechanism patterns and information flow across all models. However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix J). Moreover, as the model size increases, we found evidence that the contribution of attention heads becomes more complex. We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (See Figure 3).
Related Work
@@
Mechanistic circuit analysis has emerged as a promising approach to interpreting the internal mechanisms of Transformers (Olah et al. 2020; Nanda et al. 2023; Olsson et al. 2022; Wang et al. 2023; García-Carrasco, Maté, and Carlos Trujillo 2024; Hanna, Liu, and Variengien 2024). Existing approaches investigating internal reasoning mechanisms mainly focus on math-related tasks, elucidating the information flow for answering mathematical questions (Stolfo, Belinkov, and Sachan 2023) and examining arithmetic operations (Quirke and Barez 2024). Most pertinent to our work, Wiegreffe et al. (2024) provides a mechanistic interpretation of multiple-choice question answering, investigating attention head-level patterns. In general, our mechanistic analysis complements recent work investigating the challenges in processing reasoning arguments that contradict established beliefs and whether the reasoning in LMs stems from internalized rules or memorized content (Ando et al. 2023; Yu, Merullo, and Pavlick 2023; Wu et al. 2024; Talmor et al. 2020; Wu et al. 2024; Kassner, Krojer, and Schütze 2020; Haviv et al. 2023; Feldman 2020; Monea et al. 2024; Yu, Merullo, and Pavlick 2023; Wallat, Singh, and Anand 2020). To address this question, different mechanistic approaches have been adopted to localizing factual associations (Meng et al. 2022; Geva et al. 2023; Dai et al. 2022) or assessing conditions for generalisation (Wang et al. 2024). However, to the best of our knowledge, we are the first to investigate content-independent mechanisms for syllogistic reasoning.
Conclusion
@@
It is important to acknowledge some of the limitations in our study: (1) our analysis is conducted predominantly on the transitive property and term-specific information and does not consider the full spectrum of the reasoning dynamics that might appear in more complex scenarios. (2) Our experimentation is confined to a specific reasoning template that lacks the complexity and noise of real-world problems due to the intricate level of granularity and the computational complexity required in circuit analysis. Despite these limitations, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by auto-regressive language models and lay a solid foundation for future research in this field.
@@ Acknowledgements
This work was partially funded by the Swiss National Science Foundation (SNSF) project NeuMath: Neural Discourse Inference over Mathematical Texts (200021_204617).
Appendix
This document serves as a technical appendix to the paper titled A Mechanistic Interpretation of Syllogistic Reasoning in Auto-Regressive Language Models. It provides additional theoretical background, experimental details, and supplementary results that support the main findings presented in the original paper.
Appendix A A. Unconditionally Valid Syllogism Schemes
@@
Table 3: 15 Unconditionally valid syllogisms schemes. The table lists the syllogisms by their traditional names, moods, and figures, with formalised logical expressions on the first line and corresponding natural language examples on the second line. The minor premise (m) is presented before the major premise (M) to emphasise the transitive property inherent in these forms.
Appendix B B. Dataset Generation
B.1. Symbolic
The symbolic dataset comprises sentences where all terms in the premises are represented by abstract symbols (uppercase alphabet letters). From the set of all 26 uppercase alphabets, three-letter triples (e.g., A, B and C) are randomly sampled. For each triple, six permutated prompts and label pairs are generated following a template designed to minimise latent semantic interference among alphabet symbol tokens.
B.2. Non-symbolic
The non-symbolic datasets are constructed based on GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), a resource that provides a foundation for evaluation of sentence veracity with associated truthfulness scores (0-1). The dataset construction process involves the following steps:
  • Extraction: Select generic sentences with a truthfulness score of 1 based on GenericsKB (Bhakthavatsalam, Anastasiades, and Clark 2020), specifically those in the form A are B, using regular expression.
@@ B.2. Non-symbolic
  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI et al. 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent. This step helps us classify a belief-consistent dataset and a belief-inconsistent dataset.
@@ B.3. Data Statistics
Appendix C C. Embedding Space Analysis Details
One well-known method for analysing hidden activation in transformer language models is projecting activation into embedding space, the so-called logit lens (Nostalgebraist 2020; Elhage et al. 2021; Geva et al. 2022). In particular, we focus on an input-agnostic approach (Dar et al. 2023; Hanna, Liu, and Variengien 2024), utilising both unembedding ($\mathbf{W}_{U}$) and embedding ($\mathbf{W}_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ represents the model’s vocabulary size. In this study, we employ the OV circuit (Elhage et al. 2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand the general behaviour of how source tokens influence output logits. The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalised as $\mathbf{W}_{E}\mathbf{W}^{\mathbf{h}}_{V}\mathbf{W}_{O}^{\mathbf{h}}\mathbf{W}_{U}\in\mathbb{R}^{|V|\times|V|}$. For simplicity and in line with established conventions (Elhage et al. 2021; Dar et al. 2023), this formulation ignores layer normalisation.
Appendix D D. Circuit Ablation Method Details
@@
Appendix E E. GPT-2 Model Architecture
We provide a self-contained concise overview of the GPT-2 model architecture, highlighting the main components and their mathematical relationships. Bias terms are not presented for the simplicity. We refer the conventions of notation in Elhage et al. (2021) and (Geva et al. 2023).
Notation and Preliminaries
@@
  • $W_{Q},W_{K},W_{V}$ Query, key, and value weight matrices of attention heads.
@@
$r_{i}^{0}=x_{i}W_{E}^{\text{token}}+x_{i}W_{E}^{\text{pos}}$
@@
where $r_{i}^{l}$ represents the state of the $i$-th token at layer $l$, and $A_{i}^{l}$, $M_{i}^{l}$ are outputs from the Multi-Head Self-Attention (MHSA) and Multi-Layer Perceptron (MLP) sublayers.
Multi-Head Self-Attention (MHSA).
@@
$A_{i}^{l}=\sum_{h\in H}\left(\sigma\left(\frac{(r_{i}^{l,h}W^{h}_{Q})(r_{i}^{l,h}W^{h}_{K})^{T}}{\sqrt{d_{K}}}\right)(r_{i}^{l,h}W^{h}_{V})\right)W^{h}_{O}$
@@
Multi-Layer Perceptron (MLP).
@@
Layer Normalisation.
Before processing by MHSA and MLP, each token’s state is normalised to stabilise learning and reduce training time:
@@
Appendix F F. Path Patching Details
Path patching, an generalised version of activation patching, aims to compute direct effects among model components, rather than indirect effects diffused from intervened components (Wang et al. 2023). This method involves freezing non-targeted activations during an initial forward pass, storing the corrupted targeted activation state, and then executing a subsequent forward pass with the targeted activation state substituted by the corrupted one. This approach isolates targeted component-to-component effects, eliminating non-relevant interactions. In our implementation, we employ a noising method where clean activations are replaced with corrupted ones, resulting in negative scores indicating positive contributions from corresponding components. Our findings, illustrated in Figure 7, reveal that $\mathbf{h}_{11.10}$ strongly operates based on several heads: $\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$, and $\mathbf{h}_{7.2}$. Manual investigation of attention patterns subsequently confirmed these as induction heads (Elhage et al. 2021).

### v2

@@ 3.2 Term-Related Information Flow
Now, we use the all-term intervention method to localize mover heads that carry term-related information. In the residual stream patching results (Figure 3(d)), we observe the highest positive score at the $[p]$ position, while negative scores are most prominent at the $[m_{1}]$ and $[m_{2}]$ positions, indicating that the information residing in these positions indeed contributes to their corresponding token prediction. This observation aligns with the importance of token embedding information at their respective positions given iterative refinement on the residual stream (Simoulin and Crabbé, 2021). A closer examination of the last token position reveals that the negative effect propagates from relatively early layers (approximately layer 10 onwards), yet positive effects are from later layers (approximately layer 14 onwards), incurring a positive shift of the model’s prediction from $m$ to $p$.
@@
Given the findings that the information for prediction resides in term-specific positions, we trace which attention heads transfer the information from each term position ($[p]$, $[m_{1}]$ and $[m_{2}]$) to the last token. We call these heads mover heads as the existence of “positive or negative mover heads” that carry or suppress information from the specific token position to the last token position (Wang et al., 2023; García-Carrasco et al., 2024). To discover the source of information flow, we investigate the heads with high attention value-based patching score and identify 9 notable mover heads (see Appendix H for details of mover heads localization). By further investigation of attention patterns (Figure 3(f)–(h)), $\mathbf{h}_{14.14},\mathbf{h}_{15.14}$ and $\mathbf{h}_{18.12}$ are classified as positive copy heads attending highly to $[p]$, and $\mathbf{h}_{19.1}$ is a positive suppression head by attending to $[m_{2}]$ and $[s]$. $\mathbf{h}_{9.9}$, $\mathbf{h}_{11.1}$, $\mathbf{h}_{12.1}$, $\mathbf{h}_{17.2}$ and $\mathbf{h}_{23.10}$ are negative copy heads giving a strong attention to non-$[p]$.
@@
3.3 Summary
Overall, we found that the mechanisms to derive correct conclusion involve the following phases:
(1) Long Induction:
Early layers exhibit biases towards the wrong conclusion due to long-range repetition of the first premise “All A are B”.
(2) Duplication:
Induction heads aggregate information about duplicated middle terms in the premises.
(3) Suppression:
The model aggregates and inhibits middle-term information (i.e.,‘B’) suppressing the long induction mechanism.
(4) Mover:
Token-specific information is propagated to the last token position. The process culminates in the prediction shift from ‘B’ to the correct token ‘C’.
Therefore, these results suggest that the discovered circuit is mainly characterized by an internal error correction mechanism.
3.4 Circuit Evaluation
@@ 3.5 Circuit Transferability
We present ablation results for the two generated non-symbolic datasets: belief-consistent (Figure 5(a)) and belief-inconsistent (Figure 5(b)) (See Appendix I for the intervention results). Notably, the symbolic circuit proves necessary for both types of non-symbolic inputs, suggesting that the logic derived from symbolic syllogisms remains essential even when natural words are substituted. Regarding sufficiency, while belief-consistent data show significant performance recovery, we observe an inability to restore original performance on the belief-inconsistent set despite an increasing trend. These results indicate that belief biases encoded in different attention heads may play an important role.
@@
3.6 Generalization to Syllogistic Schemes
@@
We further aim to understand whether the symbolic circuit is specific to the AAA-1 syllogism (Barbara). To this end, we extend our experiment to encompass all 15 unconditionally valid syllogisms (see Appendix A and J for details about the syllogistic schemes and more experimental results). To evaluate the transferability to all 15 schemes, we verify three main conditions, reported in Table 2: (C1) Necessity, (C2) Sufficiency, and (C3) Positive Logit Difference. We also measure the accuracy of the completion task for each syllogism. From Table 2, we observe that the circuit is sufficient and necessary for all the syllogistic schemes in which the model achieves high downstream accuracy ($\geq$ 60%), supporting the conclusion that the circuit includes components that are crucial for the emergence of syllogistic reasoning in general. Notably, all three conditions are satisfied for the affirmative syllogisms (AII-3, IAI-3, IAI-4), for which the model achieves an accuracy above 0.6.
3.7 Generalization to Different Sizes
@@
Next, we expand our analysis to different sizes of GPT-2 (small, large and XL). Here, the intervention results show similar suppression mechanism patterns and information flow across all models. However, the residual stream of GPT-2 small in the all-term corruption setup is reversed due to its low downstream accuracy (see Appendix K for additional details). Moreover, as the model size increases, we found evidence that the contribution of attention heads becomes more complex. We hypothesize this might be caused by a stronger impact of world knowledge with increasing size, as suggested by the decrease in accuracy on the symbolic dataset and the increase on the non-symbolic set observed for the XL model (See Figure 6(a)–(b)).
3.8 Generalization to Different Models
We further extend our investigation to assess whether a similar reasoning mechanism can be identified across different LMs. Specifically, we evaluate the circuit on: Pythia with 70M, 160M, 410M, and 1B parameters (Biderman et al., 2023); Qwen 2.5 with 0.5B and 1.5B parameters (Yang et al., 2024); Llama 3.2 with 1B parameters (Grattafiori et al., 2024); and additionally, GPT-2-medium after being fine-tuned on argumentative texts for critical thinking (Betz et al., 2021).
Our results reveal distinct grouping patterns. In relatively small models—such as Pythia 70M and 160M—we observe erratic activation patterns that mirror the limitations seen in under-scaled models like GPT-2 small. In contrast, models with more than 0.5B parameters (namely, Pythia 410M, Pythia 1B, and Llama 3.2 1B) exhibit a compatible pattern in terms of information flow and suppressive mechanisms. Qwen exhibits a variant suppressive mechanism, with the suppression head operating at the final token position rather than $[p]$. Meanwhile, the fine-tuned GPT-2 medium model yields activation patterns that are nearly indistinguishable from its pre-trained counterpart, suggesting that the core reasoning circuit originates from pre-training rather than fine-tuning. Overall, our results indicate that while the core reasoning circuit generalizes effectively across sufficiently scaled models, subtle architectural nuances can lead to distinct but comparable internal dynamics (See Appendix L for the result details).
4 Related Work
@@
Mechanistic circuit analysis has emerged as a promising approach to interpreting the internal mechanisms of Transformers (Olah et al., 2020; Nanda et al., 2023; Olsson et al., 2022; Wang et al., 2023; García-Carrasco et al., 2024; Hanna et al., 2024). Existing approaches investigating internal reasoning mechanisms mainly focus on math-related tasks, elucidating the information flow for answering mathematical questions (Stolfo et al., 2023) and examining arithmetic operations (Quirke and Barez, 2024). Most pertinent to our work, Wiegreffe et al. (2024) provides a mechanistic interpretation of multiple-choice question answering, investigating attention head-level patterns. In general, our mechanistic analysis complements recent work investigating the challenges in processing reasoning arguments that contradict established beliefs and whether the reasoning in LMs stems from internalized rules or memorized content (Ando et al., 2023; Yu et al., 2023; Wu et al., 2024; Talmor et al., 2020; Wu et al., 2024; Kassner et al., 2020; Haviv et al., 2023; Feldman, 2020; Monea et al., 2024; Yu et al., 2023; Wallat et al., 2020; Eisape et al., 2024). To address this question, different mechanistic approaches have been adopted to localizing factual associations (Meng et al., 2022; Geva et al., 2023; Dai et al., 2022) or assessing conditions for generalization (Wang et al., 2024). However, to the best of our knowledge, we are the first to investigate mechanisms for syllogistic reasoning.
5 Conclusion
@@
6 Limitations
It is important to acknowledge some of the limitations of our study. First, our analysis is conducted predominantly on the transitive property and term-specific information and does not consider the full spectrum of the reasoning dynamics that might appear in more complex scenarios which are hard to model via causal intervention techniques. Second, Our experimentation is confined to specific syllogistic reasoning templates due to the intricate level of granularity and the computational complexity required in circuit analysis. Despite these limitations, however, we believe our findings could offer valuable insights into the reasoning mechanisms adopted by auto-regressive language models for reasoning along with existing biases and lay a solid foundation for future research in this field.
@@ Acknowledgements
This work was funded by the Swiss National Science Foundation (SNSF) project “NeuMath” (200021_204617), by the CRUK National Biomarker Centre, and supported by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
Appendix A Unconditionally Valid Syllogism Schemes
@@
Table 3: 15 Unconditionally valid syllogism schemes. The table lists the syllogisms by their traditional names, moods, and figures, with formalised logical expressions on the first line and corresponding natural language examples on the second line. The minor premise (m) is presented before the major premise (M) to emphasise the transitive property in AAA-1 syllogism.
Appendix B Dataset Generation
B.1 Symbolic
The symbolic dataset comprises sentences where all terms in the premises are represented by abstract symbols (uppercase alphabet letters). From the set of all 26 uppercase alphabets, three-letter triples (e.g., A, B and C) are randomly sampled. For each triple, six permutated prompts and label pairs are generated following syllogism templates designed to minimise latent semantic interference among alphabet symbol tokens.
B.2 Non-symbolic
The non-symbolic datasets are constructed based on GenericsKB (Bhakthavatsalam et al., 2020), a resource that provides a foundation for the evaluation of sentence veracity with associated truthfulness scores (0-1). The dataset construction process involves the following steps:
  • Extraction: Select generic sentences with a truthfulness score of 1 based on GenericsKB (Bhakthavatsalam et al., 2020), specifically those in the form A are B, using regular expression.
@@ B.2 Non-symbolic
  • Classification: To address issues of partial inclusion and syntactic ambiguity in constructed syllogism arguments, we employ GPT-4 (OpenAI et al., 2024) to classify whether arguments contain only truthful premises and whether the middle-terms are syntactically and semantically equivalent. This step helps us classify a belief-consistent dataset and a belief-inconsistent dataset.
@@ B.3 Data Statistics
Appendix C Embedding Space Analysis Details
One established method for analysing hidden activation in transformer language models is the logit lens, which projects activation into embedding space (Nostalgebraist, 2020; Elhage et al., 2021; Geva et al., 2022). We focus on an input-agnostic approach (Dar et al., 2023; Hanna et al., 2024), utilising both unembedding ($W_{U}$) and embedding ($W_{E}$) matrices to construct a $\mathbb{R}^{|V|\times|V|}$ matrix, where $|V|$ denotes the model’s vocabulary size. This study employs the OV circuit (Elhage et al., 2021) of attention heads, formed by attention value and output weights ($W_{V}W_{O}$), to understand how source tokens generally influence output logits. The OV circuit-based logit lens for attention head $\mathbf{h}$ is formalised as $W_{E}W^{\mathbf{h}}_{V}W_{O}^{\mathbf{h}}W_{U}\in\mathbb{R}^{|V|\times|V|}$. Following the previous works (Elhage et al., 2021; Dar et al., 2023), this formulation omits layer normalisation.
Appendix D Circuit Ablation Method Details
@@
Appendix E Phase Transition
In Figure 6(a–b), we observe a notable phase transition from small to medium models in accuracy and logit difference.
Figure 6: Accuracy and logit difference on the syllogism completion task across all different sizes of GPT-2 for three different datasets.
Appendix F GPT-2 Model Architecture
We provide a self-contained concise overview of the GPT-2 model architecture, highlighting the main components and their mathematical relationships. Bias terms are not presented for the simplicity. We refer the conventions of notation in Elhage et al. (2021) and Geva et al. (2023).
Notation
@@ Appendix F GPT-2 Model Architecture
  • $p_{i}$ One-hot encoded vector representing the $i$-th positional information.
  • $r_{i}^{l}$ Residual stream vector representing the $i$-th token at layer $l$ in the input sequence.
@@
  • $A_{i}^{l}$ Output from the Multi-Head Self-Attention sublayer
  • $M_{i}^{l}$ Output from Multi-Layer Perceptron (MLP) sublayer.
  • $W_{Q},W_{K},W_{V},W_{O}$ Query, key, value, output weight matrices of attention heads.
@@
$r_{i}^{0}=x_{i}W_{E}^{\text{token}}+p_{i}W_{E}^{\text{pos}}$
@@
Multi-Head Self-Attention Sublayer.
@@
$A_{i}^{l}=\sum_{h\in H}\left(\sigma\left(\frac{(r_{i}^{l}W^{h}_{Q})(r_{i}^{l}W^{h}_{K})^{T}}{\sqrt{d_{K}}}\right)(r_{i}^{l}W^{h}_{V})\right)W^{h}_{O}$
@@
Multi-Layer Perceptron Sublayer.
@@
Layer Normalization.
Before processing by self-attention and MLP processing, each token’s state is normalized to stabilise learning and reduce training time:
@@
Appendix G Path Patching Details
Path patching, an generalized version of activation patching, aims to compute direct effects among model components rather than indirect effects diffused from intervened components (Wang et al., 2023). This method involves freezing non-targeted activations during an initial forward pass, storing the targeted activation state in intervention, and then executing a subsequent forward pass with the targeted activation state substituted by the stored one. This approach isolates targeted component-to-component effects, eliminating non-relevant interactions. In our implementation, we employ a noising method where clean activations are replaced with corrupted ones, resulting in negative scores indicating positive contributions from corresponding components. Our findings, illustrated in Figure 7, reveal that $\mathbf{h}_{11.10}$ strongly operates based on several heads: $\mathbf{h}_{5.8}$, $\mathbf{h}_{6.1}$, $\mathbf{h}_{6.15}$, and $\mathbf{h}_{7.2}$. Manual investigation of attention patterns subsequently confirmed these as induction heads (Elhage et al., 2021).